# Informed Asymmetric Actor-Critic: Leveraging Privileged Signals Beyond Full-State Access

### v1

@@ Appendix A AUXILIARY RESULTS
Generality of state-dependent information functions.
It is worth noting that the assumption of the information function $I:\mathcal{S}\to\Delta(\mathcal{I})$ depending solely on the current state $s_{t}$ is not restrictive in the context of informed POMDPs. For instance, consider a more general setting where the information variable $i_{t}$ is sampled from a distribution conditioned on the current state, action, and next state, i.e., $i_{t}\sim\mathbb{P}(i_{t}\mid s_{t},a_{t},s_{t+1})$. In such cases, the model can be reformulated as an informed POMDP by augmenting the state space to $\widetilde{\mathcal{S}}=\mathcal{S}\times\mathcal{A}\times\mathcal{S}$, and defining a corresponding transition function $\widetilde{T}$ and reward function $\widetilde{R}$ over this augmented space. Under this formulation, the information function retains the standard informed POMDP form $\widetilde{I}:\widetilde{\mathcal{S}}\to\Delta(\mathcal{I})$, thereby demonstrating that the state-only dependency assumption is without loss of generality.
Appendix B BENCHMARK ENVIRONMENTS
@@
In this section, we describe the partially observable benchmark environments used in our experiments.
@@
The Heaven-Hell task [28, 29] is a partially observable navigation problem within a gridworld environment characterized by a corridor-like structure with a fork leading to three distinct terminal branches. Two of these branches correspond to terminal exits: one leading to a positive outcome (heaven) and the other to a negative outcome (hell). The third branch leads to a non-terminal location where the agent can interact with an oracle (referred to as a "priest") who provides information necessary to disambiguate the exits. The agent is initially unaware of which terminal corresponds to heaven.
The underlying state includes both the agent’s position and the true location of the heaven exit. As observation, however, the agent either perceives its own location or, when visiting the priest, receives an observation that reveals the location of heaven. We construct privileged partial information by adding to the agent’s location its distance to the heaven terminal using Earth Mover’s distance.
@@
The Shopping-5 environment [29] models another grid-world navigation task in which an agent must buy a forgotten item from a store. The environment is modeled as a two-dimensional gridworld of size $5\times 5$, with the item placed randomly at one of the grid cells. The agent begins at an arbitrary location and must locate and buy the item. While the agent’s position is fully observable, the item’s position is hidden and must be explicitly queried.
Hence, the full state encodes both the agent’s position and the item’s location, represented compactly as integers. Observations are similarly encoded, but are partial: at each time step, the agent observes either its own position or, upon executing a query, the position of the item. Similar to the Heaven-Hell task, we introduce a privileged partial by computing the current Earth Mover’s distance between the agent and the item.
@@
The so-called Gridverse suite [32] defines a collection of partially observable environments in which agents interact within structured gridworlds. In this work, we consider the $7\times 7$-Memory-Four-Room and $9\times 9$-Memory-Four-Room environments. While actions are encoded as categorical indices, both states and observations are structured representations comprising multiple semantically meaningful components. Importantly, these components differ between state and observation, and some are only available in the state representation. The key components are:
@@
Appendix C IMPLEMENTATION DETAILS
In this section, we detail some parts of our implementation used in the evaluation. All experiments were conducted on a cluster node equipped with 64 cores running at 3.0 GHz and 72 GB of RAM allocated per task.
C.1 Model Architectures
@@
  • For the Car-Flag and Cleaner environments, actions are represented as one-hot encodings of their respective categorical indices. As the state and observation representations provided by these environments are already flattened and structurally simple, no additional embedding is applied. The actor and critic subsequent networks adopt the same architecture used for the Heaven-Hell-3 and Shopping-5 tasks.
@@
For each environment and method, we use the hyperparameter values recommended by Baisero and Amato [10] to ensure comparability with prior work. Table 2 summarizes the actor learning rate $\alpha_{\pi}$, critic learning rate $\alpha_{\hat{V}}$, and the initial negative-entropy weight $\lambda_{0}$ selected for each environment. Additionally, the following model hyperparameters are applied across all environments: discount factor is set to $\gamma=0.99$, episodes are automatically terminated if they exceed 100 time steps; two episodes are sampled per gradient update; a frozen target network is used to stabilize critic training, with target parameters updated every 10,000 time steps; and the negative-entropy weight $\lambda$ decays linearly over 2 million time steps to a final value equal to one-tenth of $\lambda_{0}$.
Environment $\alpha_{\pi}$ $\alpha_{\hat{V}}$ $\lambda_{0}$

### v2

@@ Appendix A Auxiliary Results
Appendix B Analysis of the $\alpha$-Residual-Informativeness Test
In this section, we detail the construction of the proposed $\alpha$-residual-informativeness test and analyze its power.
B.1 Test Construction
In this subsection, we formalize the null distribution implicitly approximated by the permutation procedure used in the residual-based informativeness test. Our goal is to clarify how the proposed test relates to the conditional independence (CI) hypothesis
$\mathbb{H}_{0}^{\mathrm{CI}}:\quad G_{t}\perp\!\!\!\perp i_{t}|h_{t},a_{t}.$
Let $G_{t},i_{t},h_{t}$, and $a_{t}$ be random variables with joint distribution $p(G_{t},i_{t},h_{t},a_{t})$ and finite second moments. Define the conditional-mean residuals
$e_{G_{t}}:=G_{t}-\mathbb{E}[G_{t}|h_{t},a_{t}],\qquad e_{i_{t}}:=i_{t}-\mathbb{E}[i_{t}|h_{t},a_{t}].$
Proposition B.1.
If $\mathbb{H}_{0}^{\mathrm{CI}}$ holds, then
$e_{G_{t}}\perp\!\!\!\perp e_{i_{t}}\mid h_{t},a_{t}.$
Proof.
Under $\mathbb{H}_{0}^{\mathrm{CI}}$, the conditional joint distribution factorizes:
$p(G_{t},i_{t}|h_{t},a_{t})=p(G_{t}|h_{t},a_{t})\,p(i_{t}|h_{t},a_{t}).$
Subtracting measurable functions of $(h_{t},a_{t})$ from each variable does not affect CI given $(h_{t},a_{t})$, so
$p(e_{G_{t}},e_{i_{t}}|h_{t},a_{t})=p(e_{G_{t}}|h_{t},a_{t})\,p(e_{i_{t}}|h_{t},a_{t}).$
This concludes the proof. ∎
The converse does not hold in general [45]; thus, testing residual independence is a relaxation of CI testing.
Ideal residual null distribution.
An exact test of $\mathbb{H}_{0}^{\mathrm{res}}:e_{G_{t}}\perp\!\!\!\perp e_{i_{t}}\mid h_{t},a_{t}$ would require samples from
$p(h_{t},a_{t})\,p(e_{G_{t}}|h_{t},a_{t})\,p(e_{i_{t}}|h_{t},a_{t}),$
where the residual conditionals are induced by the joint law of $(G_{t},h_{t},a_{t})$ and $(i_{t},h_{t},a_{t})$.
While $(h_{t},a_{t},e_{i_{t}})$ can be sampled from data once $\mathbb{E}[i_{t}|h_{t},a_{t}]$ is estimated, obtaining independent samples from $p(e_{G_{t}}|h_{t},a_{t})$ is intractable because it requires sampling from the unknown return distribution conditioned on the full history $h_{t}$ and action $a_{t}$.
First-moment-matching surrogate.
To construct a tractable null distribution, we replace the original return $G_{t}$ by a surrogate variable.
Assumption B.2 (Existence of a surrogate).
There exists a random variable $\tilde{G}_{t}$ defined on the same probability space such that
  1. (i)
    $\tilde{G}_{t}\overset{d}{=}G_{t}$ (same marginal distribution),
  2. (ii)
    $\tilde{G}_{t}\perp\!\!\!\perp i_{t},h_{t},a_{t}$ (independence from the privileged signal, history, and action),
  3. (iii)
    $\mathbb{E}[\tilde{G}_{t}^{2}]<\infty$ (finite second moment).
The surrogate null is then obtained by replacing $G_{t}$ with
$\tilde{G_{t}}^{\text{null}}:=\tilde{G}_{t}-\mathbb{E}[\tilde{G}_{t}]+\mathbb{E}[G_{t}|h_{t},a_{t}].$ (1)
The corresponding surrogate null residual is
$e_{\tilde{G}_{t}}^{\text{null}}:=\tilde{G}^{\text{null}}_{t}-\mathbb{E}[G_{t}|h_{t},a_{t}]=\tilde{G}_{t}-\mathbb{E}[\tilde{G}_{t}].$
Proposition B.3 (Properties of the surrogate residual).
For all $(h_{t},a_{t})$,
$\mathbb{E}[e_{\tilde{G}_{t}}^{\text{null}}|h_{t},a_{t}]=0,\qquad\text{and}\qquad e_{\tilde{G}_{t}}^{\text{null}}\perp\!\!\!\perp i_{t}|h_{t},a_{t}.$
Proof.
First, we prove zero conditional mean. By definition,
$e_{\tilde{G}_{t}}^{\text{null}}=\tilde{G}_{t}-\mathbb{E}[\tilde{G}_{t}],$
where $\mathbb{E}[\tilde{G}_{t}]$ is a constant. Taking the conditional expectation given $(h_{t},a_{t})$, yields
$\mathbb{E}[e_{\tilde{G}_{t}}^{\text{null}}|h_{t},a_{t}]=\mathbb{E}[\tilde{G}_{t}|h_{t},a_{t}]-\mathbb{E}[\tilde{G}_{t}].$
Since $\tilde{G}_{t}$ is independent of $(h_{t},a_{t})$ by Assumption B.2(ii),
$\mathbb{E}[\tilde{G}_{t}|h_{t},a_{t}]=\mathbb{E}[\tilde{G}_{t}].$
Therefore,
$\mathbb{E}[e_{\tilde{G}_{t}}^{\text{null}}|h_{t},a_{t}]=0.$
Second, we show conditional independence from $i_{t}$. Again using
$e_{\tilde{G}_{t}}^{\text{null}}=\tilde{G}_{t}-\mathbb{E}[\tilde{G}_{t}],$
note that subtracting a constant does not affect independence relations. Thus, it suffices to show
$\tilde{G}_{t}\perp\!\!\!\perp i_{t}|h_{t},a_{t}.$
For any measurable sets $\mathcal{X}$, $\mathcal{Y}$,
$p(\tilde{G}_{t}\in\mathcal{X},i_{t}\in\mathcal{Y}|h_{t},a_{t})=\frac{p(\tilde{G}_{t}\in\mathcal{X},i_{t}\in\mathcal{Y},h_{t},a_{t})}{p(h_{t},a_{t})}.$ (2)
Since $\tilde{G}_{t}$ is independent of $(h_{t},a_{t},i_{t})$,
$p(\tilde{G}_{t}\in\mathcal{X},i_{t}\in\mathcal{Y},h_{t},a_{t})=p(\tilde{G}_{t}\in\mathcal{X})\ p(i_{t}\in\mathcal{Y},h_{t},a_{t}).$
Substituting into Equation 2 gives
$p(\tilde{G}_{t}\in\mathcal{X},i_{t}\in\mathcal{Y}|h_{t},a_{t})$ $=p(\tilde{G}_{t}\in\mathcal{X})\ \frac{p(i_{t}\in\mathcal{Y},h_{t},a_{t})}{p(h_{t},a_{t})}$
$=p(\tilde{G}_{t}\in\mathcal{X})\ p(i_{t}\in\mathcal{Y}|h_{t},a_{t}).$
Finally, because $\tilde{G}_{t}\perp\!\!\!\perp h_{t},a_{t}$, we have
$p(\tilde{G}_{t}\in\mathcal{X})=p(\tilde{G}_{t}\in\mathcal{X}|h_{t},a_{t}).$
Therefore,
$p(\tilde{G}_{t}\in\mathcal{X},i_{t}\in\mathcal{Y}|h_{t},a_{t})=p(\tilde{G}_{t}\in\mathcal{X}|h_{t},a_{t})\ p(i_{t}\in\mathcal{Y}|h_{t},a_{t}).$
which establishes
$\tilde{G}_{t}\perp\!\!\!\perp i_{t}|h_{t},a_{t}.$
Hence,
$e_{\tilde{G}_{t}}^{\text{null}}\perp\!\!\!\perp i_{t}|h_{t},a_{t}.$
This concludes the proof. ∎
Thus, the surrogate residual distribution
$p(h_{t},a_{t})\,p(e_{\tilde{G}_{t}}^{\text{null}}|h_{t},a_{t})\,p(e_{i_{t}}|h_{t},a_{t})$
matches the ideal residual-based null in the conditional mean of the residual while enforcing CI.
Therefore, the permutation test operates under a first-moment-matching surrogate null rather than the ideal CI null. Nevertheless, the test targets violations of residual independence and allows for identifying privileged signals that provide additional predictive information about returns. Moreover, the proposed test has the following interpretation:
  • If CI holds, then residuals are conditionally independent given $(h_{t},a_{t})$, and the test does not reject $\mathbb{H}_{0}^{\mathrm{res}}$ asymptotically under standard regularity conditions.
  • If the test rejects $\mathbb{H}_{0}^{\mathrm{res}}$, then $i_{t}$ contains predictive information about the conditional mean of $G_{t}$ beyond what is captured by $\mathbb{E}[G_{t}|h_{t},a_{t}]$.
  • The test does not guarantee full CI, only the absence of residual predictive information.
However, this notion of informativeness is in line with the objective of value-function learning, where conditional means, not full conditional distributions, determine optimal predictions.
B.2 Power Analysis of the $\alpha$-Residual-Informativeness Test
This subsection studies how errors in the conditional expectation estimators affect the population dependence measure underlying the residual-informativeness test.
Regression error and residual perturbation.
Let $m_{G}(h_{t},a_{t})=\mathbb{E}[G_{t}|h_{t},a_{t}]$ and $m_{i}(h_{t},a_{t})=\mathbb{E}[i_{t}|h_{t},a_{t}]$ denote the true conditional expectations, and let $\hat{m}_{G}(h_{t},a_{t}),\hat{m}_{i}(h_{t},a_{t})$ be learned estimators. Define the regression errors
$\Delta_{G_{t}}$ $=m_{G}(h_{t},a_{t})-\hat{m}_{G}(h_{t},a_{t}),$ (3)
$\Delta_{i_{t}}$ $=m_{i}(h_{t},a_{t})-\hat{m}_{i}(h_{t},a_{t}).$ (4)
We assume the regression errors have finite second moments, i.e., $\mathbb{E}[\Delta^{2}_{G_{t}}]<\infty$, and $\mathbb{E}[\Delta^{2}_{i_{t}}]<\infty$.
The population residuals are
$e_{G_{t}}=G_{t}-m_{G}(h_{t},a_{t}),\qquad e_{i_{t}}=i_{t}-m_{i}(h_{t},a_{t}),$
while the estimated residuals are
$\hat{e}_{G_{t}}=G_{t}-\hat{m}_{G}(h_{t},a_{t})=e_{G_{t}}+\Delta_{G_{t}},\qquad\hat{e}_{i_{t}}=i_{t}-\hat{m}_{i}(h_{t},a_{t})=e_{i_{t}}+\Delta_{i_{t}}.$
Throughout this section, we assume sample splitting, following the procedure described in Algorithm 1, so that regression estimation is independent of the dependence test.
Let $\rho(X,Y)$ denote a population dependence measure between square-integrable random variables. We impose the following stability condition.
Assumption B.4 (Lipschitz continuity in $L^{2}$).
There exists a constant $L_{\rho}>0$ such that for all $X,X^{\prime},Y,Y^{\prime}$ with finite second moments,
$\big|\rho(X,Y)-\rho(X^{\prime},Y^{\prime})\big|\leq L_{\rho}\big(\|X-X^{\prime}\|_{2}+\|Y-Y^{\prime}\|_{2}\big),$
where $\|Z\|_{2}:=\sqrt{\mathbb{E}[Z^{2}]}$.
This assumption is satisfied by many kernel- and distance-based dependence measures under additional regularity conditions (e.g., bounded Lipschitz kernels and finite second moments).
Applying Assumption B.4 with
$(X,Y)=(e_{G_{t}},e_{i_{t}}),\qquad(X^{\prime},Y^{\prime})=(\hat{e}_{G_{t}},\hat{e}_{i_{t}}),$
yields
$\big|\rho(e_{G_{t}},e_{i_{t}})-\rho(\hat{e}_{G_{t}},\hat{e}_{i_{t}})\big|\leq L_{\rho}\big(\|\Delta_{G_{t}}\|_{2}+\|\Delta_{i_{t}}\|_{2}\big).$ (5)
Thus, regression error perturbs the population dependence measure by at most an additive margin proportional to the $L^{2}$ regression errors.
Implications for the detectable signal strength.
Let $q_{\alpha}$ denote the rejection threshold of a level-$\alpha$ test based on an estimator of the dependence measure. Ignoring sampling variability for the moment and focusing on the underlying population quantities, the Lipschitz bound in Equation 5 implies
$\big|\rho(e_{G_{t}},e_{i_{t}})-\rho(\hat{e}_{G_{t}},\hat{e}_{i_{t}})\big|\leq\Delta_{\text{reg}},\qquad\Delta_{\text{reg}}=L_{\rho}\big(\|\Delta_{G_{t}}\|_{2}+\|\Delta_{i_{t}}\|_{2}\big).$
This yields two implications:
  1. 1.
    If the estimated dependence exceeds the rejection threshold, then the true residual dependence cannot be much smaller:
    $\rho(\hat{e}_{G_{t}},\hat{e}_{i_{t}})>q_{\alpha}\;\Rightarrow\;\rho(e_{G_{t}},e_{i_{t}})>q_{\alpha}-\Delta_{\text{reg}}.$
    Thus, false rejections due purely to regression error are controlled up to an additive tolerance $\Delta_{\text{reg}}$.
  2. 2.
    Conversely, if the true residual dependence is sufficiently strong, then it remains detectable despite regression error perturbation:
    $\rho(e_{G_{t}},e_{i_{t}})>q_{\alpha}+\Delta_{\text{reg}}\;\Rightarrow\;\rho(\hat{e}_{G_{t}},\hat{e}_{i_{t}})>q_{\alpha}.$
In other words, regression error effectively increases the minimum population dependence required for reliable detection by an additive margin $\Delta_{\text{reg}}$. Alternatives with true residual dependence exceeding $q_{\alpha}+\Delta_{\text{reg}}$ remain well separated from the null, while alternatives within a $\Delta_{\text{reg}}$-neighborhood of the threshold may become indistinguishable due to estimation error.
Note that the preceding argument describes population-level perturbation. In finite samples, additional variability in the test statistic introduces further uncertainty.
HSIC with Gaussian RBF kernels as dependency measure.
We now consider $\rho$ given by the population Hilbert-Schmidt Independence Criterion (HSIC) [19] with Gaussian RBF kernels
$k(x,x^{\prime})=\exp\!\left(-\frac{|x-x^{\prime}|^{2}}{2\sigma_{x}^{2}}\right),\qquad l(y,y^{\prime})=\exp\!\left(-\frac{|y-y^{\prime}|^{2}}{2\sigma_{y}^{2}}\right).$
Gaussian RBF kernels are bounded and smooth, and are Lipschitz on bounded domains. Under finite second moments of the residuals (or sub-Gaussian tails), HSIC satisfies a bound of the form in Assumption B.4 under $L^{2}$ perturbations, with a constant that can be approximated by
$L_{\rho}\approx C\Big(\frac{1}{\sigma_{x}^{2}}+\frac{1}{\sigma_{y}^{2}}\Big),$
where $C$ depends on the kernel bounds and residual moments.
In practice, $\sigma_{x}$ and $\sigma_{y}$ are often chosen by the median-pairwise-distance heuristic. This makes $L_{\rho}$ data-dependent. Typically, residuals with larger typical pairwise distances lead to larger bandwidths and hence smaller Lipschitz constants, which can reduce the sensitivity of HSIC to small regression-induced perturbations. This effect is heuristic rather than guaranteed, but commonly observed.
If the regression estimators are $L^{2}$-consistent, i.e.,
$\|\Delta_{G_{t}}\|_{2}\to 0,\qquad\|\Delta_{i_{t}}\|_{2}\to 0,$
as the regression sample size increases, then $\Delta_{\text{reg}}\to 0$. If these errors vanish at a rate sufficiently fast relative to the sampling variability of the HSIC estimator, then the residual-informativeness test based on estimated residuals inherits the asymptotic power of the oracle test. In this sense, regression error degrades power in a controlled and continuous manner and disappears in the large-sample limit under standard consistency conditions.
Appendix C Environments
@@
This section details the partially observable environments used in our experiments.
C.1 Benchmark Environments
In this subsection, we present the partially observable benchmark environments used to assess the learning performance of the proposed informed asymmetric actor-critic method.
@@ C.1 Benchmark Environments
The Heaven-Hell task [17, 7] is a partially observable navigation problem within a grid-world environment characterized by a corridor-like structure with a fork leading to three distinct terminal branches. Two of these branches correspond to terminal exits: one leading to a positive outcome (heaven) and the other to a negative outcome (hell). The third branch leads to a non-terminal location where the agent can interact with an oracle (referred to as a "priest") who provides information necessary to disambiguate the exits. The agent is initially unaware of which terminal corresponds to heaven.
The underlying state includes both the agent’s position and the true location of the heaven exit. As observation, however, the agent either perceives its own location or, when visiting the priest, receives an observation that reveals the location of heaven. We construct privileged partial information by adding to the agent’s location its distance to the heaven terminal using Manhattan distance.
@@
The Shopping-5 environment [7] models another grid-world navigation task in which an agent must buy a forgotten item from a store. The environment is modeled as a two-dimensional grid-world of size $5\times 5$, with the item placed randomly at one of the grid cells. The agent begins at an arbitrary location and must locate and buy the item. While the agent’s position is fully observable, the item’s position is hidden and must be explicitly queried.
Hence, the full state encodes both the agent’s position and the item’s location, represented compactly as integers. Observations are similarly encoded, but are partial: at each time step, the agent observes either its own position or, upon executing a query, the position of the item. Similar to the Heaven-Hell task, we introduce a privileged partial by computing the current Manhattan distance between the agent and the item.
@@
The so-called Gridverse suite [6] defines a collection of partially observable environments in which agents interact within structured grid-worlds. In this work, we consider the $7\times 7$-Memory-Four-Room and $9\times 9$-Memory-Four-Room environments. While actions are encoded as categorical indices, both states and observations are structured representations comprising multiple semantically meaningful components. Importantly, these components differ between state and observation, and some are only available in the state representation. The key components are:
@@
C.2 Synthetic informed POMDPs
We generate a distribution of synthetic informed POMDP instances with a finite state space of size $|\mathcal{S}|$, a discrete action space of size $|\mathcal{A}|$, and continuous observation and information spaces. Following the methodology of François-Lavet et al. [15], transition probabilities are randomly assigned by setting each $(s_{t},a_{t},s_{t+1})$-entry to zero with probability $0.75$, and sampling uniformly from $[0,1]$ otherwise. To ensure valid transitions, we assign a non-zero probability to a randomly chosen next state whenever all transitions from a given state-action pair are initially zero. We then normalize the probabilities so that they sum to one. Each state is associated with a Gaussian feature vector $s_{t}\in\mathbb{R}^{d_{s}}$ with $d_{s}\in\mathbb{Z}_{>0}$:
$s_{t}\sim\mathcal{N}(\mathbf{0},\sigma_{s}^{2}\mathbf{I}_{d_{s}}),$
where $\mathbf{0}$ is the $d_{s}$-dimensional zero vector and $\mathbf{I}_{d_{s}}$ is the $d_{s}\times d_{s}$ identity matrix. Rewards are linear functions of the state features, with reward weights $w_{r}\in\mathbb{R}^{d_{s}}$ sampled uniformly from [1,1][-1,1] at initialization.
We first construct the privileged information $i_{t}\in\mathbb{R}^{d_{i}}$ by selecting a subset of $1\leq d_{i}\leq d_{s}$ state features using a binary masking vector $m_{i}\in\{0,1\}^{d_{s}}$ and applying a selection matrix $W_{i}\in\{0,1\}^{d_{i}\times d_{s}}$:
$i_{t}=W_{i}(m_{i}\odot s_{t})$
where $\odot$ denotes element-wise multiplication. Observations $o_{t}\in\mathbb{R}^{d_{o}}$ are generated by masking a subset of $1\leq d_{o}\leq d_{i}$ features from $i_{t}$ using a binary masking vector $m_{o}\in\{0,1\}^{d_{i}}$, applying a selection matrix $W_{o}\in\{0,1\}^{d_{o}\times d_{i}}$ to the masked privileged signal, and adding Gaussian noise:
$o_{t}=W_{o}(m_{o}\odot i_{t})+\beta_{o}\epsilon_{o},\qquad\epsilon_{o}\sim\mathcal{N}(\mathbf{0},\sigma^{2}_{o}\mathbf{I}_{d_{o}}),$
where $\beta_{o}\geq 0$ modulates the observation noise.
Appendix D Implementation Details
In this section, we detail parts of our implementation used in the evaluation. All experiments were conducted on a cluster node equipped with 64 cores running at 3.0 GHz and 72 GB of RAM allocated per task.
D.1 Model Architectures
@@
  • For the Car-Flag and Cleaner environments, actions are represented as one-hot encodings of their respective categorical indices. As the state and observation representations provided by these environments are already flattened and structurally simple, no additional embedding is applied. The actor’s and critic’s subsequent networks adopt the same architecture used for the Heaven-Hell-3 and Shopping-5 tasks.
@@
For each environment and method, we use the hyperparameter values recommended by Baisero and Amato [4] to ensure comparability with prior work. Table 2 summarizes the actor learning rate $\eta_{\pi}$, critic learning rate $\eta_{\hat{V}}$, and the initial negative-entropy weight $\lambda_{0}$ selected for each environment. Additionally, the following model hyperparameters are applied across all environments: discount factor is set to $\gamma=0.99$, episodes are automatically terminated if they exceed 100 time steps; two episodes are sampled per gradient update; a frozen target network is used to stabilize critic training, with target parameters updated every 10,000 time steps; and the negative-entropy weight $\lambda$ decays linearly over 2 million time steps to a final value equal to one-tenth of $\lambda_{0}$.
Table 2: Hyperparameters for the benchmark environments.
Environment $\eta_{\pi}$ $\eta_{\hat{V}}$ $\lambda_{0}$